I have a rectangle with 2 sides equaling h and the other two equaling h+4....the area is then 4h+8, but I have to find the area when a 3 inch frame is added. How do I do that?
Found 2 solutions by checkley79, stanbon:
Answer by checkley79(3341)

2(H+3)+2(H+4+3)
2H+6+2H+2*7
2H+6+2H+14
4H+20
4(H+5) ANS.

Answer by stanbon(75887)

I have a rectangle with 2 sides equaling h and the other two equaling h+4....the area is then h^2+4h, But I have to find the area when a 3 inch frame is added. How do I do that?
----
Draw the picture.
You have rectangle with base = h+4 and height = h.
It's area is h(h+4) = h^2+4h
------
Draw a 3" frame around the rectangle.
---
The new area (including the inside rectangle) has base (h+4)+6
and height = h + 6
------
The new total area is (h+10)(h+6) = h^2+16h+60
===========
Area of frame = (total area)-(original area)
= h^2+16h+60 - (h^2+4h)
= 12h + 60
=================
Cheers,
Stan H.
=================